OpenStudy (anonymous):

Divide and simplify completely. 8x - 8 x + 1 24 ______ * __________ / ______ x^2 - 1 6x^2 - 12x x^2 - 4

OpenStudy (espex):

Any time you divide fractions you are able to multiply by the inverse. I would suggest starting with your last term and invert so you can multiply to give you something like \[\frac{8x - 8}{x^2 - 1}*\frac{x + 1 }{6x^2 - 12x}*\frac{x^2 - 4}{ 24}\] Then factor your terms so that you can cancel some terms before you start multiplying.

OpenStudy (anonymous):

can you tell me what it is plzz

OpenStudy (espex):

No, you wouldn't learn anything that way and the code of conduct for OS dictates that I help others learn and prohibits me from doing the work and providing an answer.

OpenStudy (anonymous):